OpenStudy (anonymous):

Taylor's Formula Write taylor's formula for F(x,y)= sin(x)sin(y) using a=0, b=0, and n=2

OpenStudy (anonymous):

umadbro?

OpenStudy (anonymous):

He mad son.

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

what's this from? university maths?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

which year?

OpenStudy (anonymous):

which year am i?

OpenStudy (anonymous):

yes are you an undergrad?

OpenStudy (anonymous):

yes. i'm 2nd year.

OpenStudy (kinggeorge):

n=2, so we only have go up to the second degree correct?

OpenStudy (anonymous):

correct.

OpenStudy (kinggeorge):

So then the first term would merely be f(0, 0)=0

OpenStudy (kinggeorge):

Now we need to make a list of 5 partial derivatives\[\large f_{x}\]\[\large f_{y}\]\[\large f_{xx}\]\[\large f_{yy}\]\[\large f_{xy}=f_{yx}\]

OpenStudy (anonymous):

\[\sin(h)\sin(k) = hk - (1/6)h(h^2 + 3k^2)\cos \theta h \sin \theta k - (1/6)k (3h^2 + k^2) \sin \theta h \cos \theta k\] that's the answer.

OpenStudy (kinggeorge):

From now on, when I type one of the partial derivatives, assume I'm evaluating it at the point (0, 0). The formula you now need to calculate is given by\[\large f(0, 0) +f_x+f_y+f_{xy}\cdot ( xy)+{f_{xx} \over 2} \cdot(x^2)+{f_{yy} \over 2}\cdot(x^2)\]

OpenStudy (kinggeorge):

Since your function is originally, \(F(x)=\sin(x)\sin(y)\), the only nonzero term will be the \(f_{xy}\) term. All other terms will contain a \(\sin(0)=0\) term.

OpenStudy (kinggeorge):

Hence, the taylor polynomial of degree two approximating \(F(x, y)=\sin(x)\sin(y)\) is equal to \(G(x, y)=xy\)

OpenStudy (kinggeorge):

I honestly have no idea where the answer you were given comes from

OpenStudy (anonymous):

Comes from the back of the book.

OpenStudy (anonymous):

What you were doing is the taylor series i believe. We gotta use the taylor's formula.

OpenStudy (anonymous):

Apparently, there's a difference... I have no idea.

OpenStudy (kinggeorge):

As far as I can tell, there's no difference. But if there is, I've never heard of the Taylor Formula. btw, in the answer you gave me, is there some stuff cut off at the edge? Or is that all of it?

OpenStudy (anonymous):

there should be a cos(theta)k at the end of it.

OpenStudy (kinggeorge):

If that is the correct answer, I'm really confused where the theta's came from. Also, is the question given in x and y and the answer in h and k?

OpenStudy (anonymous):

Hahaha, then you're in the same boat as me. Eliassaab is an ex math professor, hopefully he'll be able to help out :)

OpenStudy (kinggeorge):

I've got to sleep now, so good luck with this problem.

OpenStudy (anonymous):

Thanks KG! I'll let you know how it goes :)

OpenStudy (inkyvoyd):

I was going to do this, but then I realized that this is probably harder math than I know. Is this after calculus, or calculus itself?

OpenStudy (anonymous):

This is calculus in its most theoretical form.

OpenStudy (inkyvoyd):

Oh,I see why I don't get it. THere are partial derivatives in here.

OpenStudy (inkyvoyd):

(x,y) lol

OpenStudy (inkyvoyd):

If you really need help, go to r/cheatatmathhomework, or math stack exchange

OpenStudy (anonymous):

Noooo idea what you are talking about.

OpenStudy (inkyvoyd):

Reddit. Or mathstackexchange.

OpenStudy (inkyvoyd):

www.reddit.com/r/cheatatmathhomework/ math.stackexchange.com/

OpenStudy (anonymous):

good luck ;)

OpenStudy (anonymous):

Awww you suck reb.

OpenStudy (anonymous):

here is tour ans.. @kevo

OpenStudy (anonymous):

I know..but I'm so good at it ;) hehe

OpenStudy (anonymous):

so the formula you want to apply reads F(a+h,b+k)=∑ i,j=0 n−1 1 i!j! ⋅∂ i+j F ∂x i ∂y j (a,b)h i k j +∑ i+j=n 1 i!j! ⋅∂ 2 F ∂x i ∂y j (a+θh,b+θk)h i k j and tells you how F behaves in a neighbourhood of (a,b) . The corresponding theorem says the for each (h,k) there is some θ∈[0,1] such that this formula holds. Now plug in your F , your n , and your (a,b) .

OpenStudy (anonymous):

then you should get your answer

OpenStudy (anonymous):

\[F(0,0) +x F^{(1,0)}(0,0)+y F^{(0,1)}(0,0)+\\ \frac{1}{2} \left(x^2 F^{(2,0)}(0,0)+2x y F^{(1,1)}(0,0)+y^2 F^{(0,2)}(0,0)\right)= x y \] All the terms are zero at (0,0) because of sin(x) except the coeeficient of x y. That is why the answer is x y